Given the probability density function
f(x)=1/4
over the interval [3,7], find the expected value, the mean, the variance and the standard deviation.
Expected value:

Mean:

Variance:

Standard Deviation:

Answer by Boreal(15235)

This is a uniform function.
mean or E(X)= (b+a)/2=5
variance is (b-a)^2/12=16/12=4/3
sd is sqrt(V)=(4/3) sqrt (3)